# Enhanced Radio Emission Between a Galaxy Cluster Pair

### v1

@@ 4 Discussion
The power of radio halos is known to strongly correlate with cluster mass. This – relation has been the focus of several statistical studies, mostly relying on SZ-derived cluster masses from Planck (e.g. R. Cassano et al., 2013; S. W. Duchesne et al., 2021a; V. Cuciti et al., 2021, 2023; R. J. van Weeren et al., 2021). As G279+39 is a Planck cluster, we can assess where it lies in the – plane by using the most recent relation derived at a frequency close to our MeerKAT 822 MHz observation; that is, the relation at 1.4 GHz obtained by V. Cuciti et al. (2021). The extrapolated power of G279+39 at 1.4 GHz is W Hz-1, placing it a factor of 2 below the best-fit relation222We refer to the “radio halos only” BCES Y—X fit of the statistical sample., in the region mostly occupied by upper limits and ultra-steep spectrum radio halos. It is worth noting that the mass of G279+39 is close to the mass cut of of the sample considered by V. Cuciti et al. (2021), therefore it lies near the boundary where the relation becomes less constrained. Moreover, the mass estimate from Planck for cluster pairs at moderate redshift is limited by the low angular resolution of the instrument, which prevents it from resolving the individual SZ components of closely spaced systems, potentially biasing the mass measurement (see A. Botteon et al., 2018, for a similar argument on A1758N-S). In these cases, the Planck mass is likely more representative of the most massive component, that in G279+39 is the eastern subcluster, whose X-ray-derived mass of (S. Döner et al., 2025) is indeed consistent within uncertainties with that reported by Planck. Overall, it seems reasonable to conclude that G279+39 falls below the – relation of radio halos. We note that, in the turbulent reacceleration scenario, underluminous radio halos are generally expected in the early or late stages of cluster mergers (J. M. Donnert et al., 2013).
@@
Lastly, we comment on the non-detection of diffuse emission in the uGMRT data (Figs. 6 and 7), showing that the emission recovered by MeerKAT is too faint to obtain a meaningful limit on the spectrum using the available uGMRT observation. As an optimistic estimation, we calculate the upper limit on the flux density at 400 MHz as , where is the ratio between the area covered by the emission to that of the beam. We adopt the level as the threshold, as it represents the limit above which the flux density of an extended radio source can be considered reliable (e.g. A. Botteon et al., 2020b). We assume that the emission covers the same area covered in the MeerKAT image, adopting an ellipse with kpc2. By considering the resolution and noise of the uGMRT image of Fig. 7 (left panel), the upper limit on the 400 MHz flux density becomes mJy. This translates to an upper limit on the spectral index of , depending on whether the integrated flux density at the or level in MeerKAT is used. A more stringent limit at the level on the 400 MHz flux density would instead imply . As anticipated, these values, which were obtained using an optimistic approach, are not useful to derive meaningful insights on the spectral index (and thus on the origin) of the diffuse emission. A more rigorous method would require to inject fake radio emission in the visibility data to estimate the upper limit, taking fully into account the flux density losses due to the missing short spacings and sparse uv-coverage of the interferometric data (see e.g. M. Johnston-Hollitt & L. Pratley, 2017; L. T. George et al., 2021; L. Bruno et al., 2023). We do not pursue this approach here as it would lead to even looser constraints. The reader is informed that the reported upper limits, despite already being relatively weak, were deliberately derived under favorable assumptions.
@@ 5 Conclusions
We reported on MeerKAT 822 MHz and uGMRT 400 MHz observations of PSZ2 G279.79+39.09 (G279+39), a poorly studied galaxy cluster pair at , likely observed shortly after pericenter passage. In X-rays, the two clusters, which are separated by a projected distance of 1.3 Mpc, are connected by a bridge of thermal gas. Our MeerKAT radio images revealed the presence of diffuse emission in the pair, while the uGMRT ones resulted in a non-detection due to the lower sensitivity. The emission is stretched along the east-west direction, with a projected size of 1500 kpc 800 kpc and a radio power at 822 MHz integrated above 3 (2) of W Hz-1 ( W Hz-1).
@@
From the flux density profile extracted along the cluster pair, we found indication that the emission in-between the two subclusters has been enhanced during the encounter. We discussed the properties of the diffuse radio emission, considered as a whole, in relation to those of radio halos and the radio bridges in A399-A401 and A1758N-S, while acknowledging the uncertainties related to the unknown 3D separation, geometry and spectral index of the emission. In principle, the estimated emissivity and radio power would be in line with that expected by a diffuse emission generated by turbulence in an early stage of the merger. However, further observations are needed to confirm the assumptions of our analysis. On the one hand, new radio data are crucial to obtain a reliable measurement of the spectral index, as the upper limits derived from the available uGMRT data are not sufficiently constraining. On the other hand, optical and high-resolution SZ observations would help to clarify the system geometry, enabling a more thorough investigation of the physical properties of the emission.
@@
To conclude, the newly discovered radio emission in G279+39 highlights the presence of nonthermal phenomena in-between a cluster pair, likely driven by turbulence injected shortly after pericenter passage. This work contributes to the growing effort to investigate the properties of diffuse radio emission in the Universe beyond the denser ICM regions typically probed by observations by targeting systems with multiple cluster components. This field is likely to flourish with the advent of next-generation facilities, with the SKA leading the way.

### v2

@@ 4 Discussion
The power of radio halos is known to strongly correlate with cluster mass. This – relation has been the focus of several statistical studies, mostly relying on SZ-derived cluster masses from Planck (e.g. Cassano et al., 2013; Duchesne et al., 2021a; Cuciti et al., 2021, 2023; van Weeren et al., 2021). As G279+39 is a Planck cluster, we can assess where it lies in the – plane by using the most recent relation derived at a frequency close to our MeerKAT 822 MHz observation; that is, the relation at 1.4 GHz obtained by Cuciti et al. (2021). The extrapolated power of G279+39 at 1.4 GHz is W Hz-1, placing it a factor of 2 below the best-fit relation222We refer to the “radio halos only” BCES Y—X fit of the statistical sample., in the region mostly occupied by upper limits and ultra-steep spectrum radio halos. It is worth noting that the mass of G279+39 is close to the mass cut of of the sample considered by Cuciti et al. (2021), therefore it lies near the boundary where the relation becomes less constrained. Moreover, the mass estimate from Planck for cluster pairs at moderate redshift is limited by the low angular resolution of the instrument, which prevents it from resolving the individual SZ components of closely spaced systems, potentially biasing the mass measurement (see Botteon et al., 2018, for a similar argument on A1758N-S). In these cases, the Planck mass is likely more representative of the most massive component, that in G279+39 is the eastern subcluster. Overall, it seems reasonable to conclude that G279+39 falls below the – relation of radio halos. We note that, in the turbulent reacceleration scenario, underluminous radio halos are generally expected in the early or late stages of cluster mergers (Donnert et al., 2013).
@@
Lastly, we comment on the non-detection of diffuse emission in the uGMRT data (Figs. 6 and 7), showing that the emission recovered by MeerKAT is too faint to obtain a meaningful limit on the spectrum using the available uGMRT observation. As an optimistic estimation, we calculate the upper limit on the flux density at 400 MHz as , where is the ratio between the area covered by the emission to that of the beam. We adopt the level as the threshold, as it represents the limit above which the flux density of an extended radio source can be considered reliable (e.g. Botteon et al., 2020a). We assume that the emission covers the same area covered in the MeerKAT image, adopting an ellipse with kpc2. By considering the resolution and noise of the uGMRT image of Fig. 7 (left panel), the upper limit on the 400 MHz flux density becomes mJy. This translates to an upper limit on the spectral index of , depending on whether the integrated flux density at the or level in MeerKAT is used. A more stringent limit at the level on the 400 MHz flux density would instead imply . As anticipated, these values, which were obtained using an optimistic approach, are not useful to derive meaningful insights on the spectral index (and thus on the origin) of the diffuse emission. A more rigorous method would require to inject fake radio emission in the visibility data to estimate the upper limit, taking fully into account the flux density losses due to the missing short spacings and sparse uv-coverage of the interferometric data (see e.g. Johnston-Hollitt & Pratley, 2017; George et al., 2021; Bruno et al., 2023). We do not pursue this approach here as it would lead to even looser constraints. The reader is informed that the reported upper limits, despite already being relatively weak, were deliberately derived under favorable assumptions.
@@ 5 Conclusions
We reported on MeerKAT 822 MHz and uGMRT 400 MHz observations of PSZ2 G279.79+39.09 (G279+39), a poorly studied galaxy cluster pair at , likely observed close (either before or after) pericenter passage. In X-rays, the two clusters, which are separated by a projected distance of 1.3 Mpc, are connected by a bridge of thermal gas. Our MeerKAT radio images revealed the presence of diffuse emission in the pair, while the uGMRT ones resulted in a non-detection due to the lower sensitivity. The emission is stretched along the east-west direction, with a projected size of 1500 kpc 800 kpc and a radio power at 822 MHz integrated above 3 (2) of W Hz-1 ( W Hz-1).
@@
From the flux density profile extracted along the cluster pair, we found indication that the emission in-between the two subclusters has been enhanced during the encounter. We discussed the properties of the diffuse radio emission, considered as a whole, in relation to those of radio halos and the radio bridges in A399-A401 and A1758N-S, while acknowledging the uncertainties related to the unknown 3D separation, geometry and spectral index of the emission. In principle, the estimated emissivity and radio power would be in line with that expected by a diffuse emission generated by turbulence in an early phase before or after the collision. However, further observations are needed to confirm the assumptions of our analysis. On the one hand, new radio data are crucial to obtain a reliable measurement of the spectral index, as the upper limits derived from the available uGMRT data are not sufficiently constraining. On the other hand, optical and high-resolution SZ observations would help to clarify the system geometry, enabling a more thorough investigation of the physical properties of the emission.
@@
To conclude, the newly discovered radio emission in G279+39 highlights the presence of nonthermal phenomena in-between a cluster pair, likely driven by turbulence injected during the cluster-cluster encounter. This work contributes to the growing effort to investigate the properties of diffuse radio emission in the Universe beyond the denser ICM regions typically probed by observations by targeting systems with multiple cluster components. This field is likely to flourish with the advent of next-generation facilities, with the SKA leading the way.
Acknowledgements.
We thank the referee for the thoughtful comments on the merger scenario and Gianfranco Brunetti, Kosuke Nishiwaki, and John ZuHone for valuable discussion. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. We thank the staff of the GMRT for support. GMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research. This research made use of the LOFAR-IT computing infrastructure supported and operated by INAF, including the resources within the PLEIADI special “LOFAR” project by USC-C of INAF, and by the Physics Dept. of Turin University (under the agreement with Consorzio Interuniversitario per la Fisica Spaziale) at the C3S Supercomputing Centre, Italy. This work made use of the following python packages: APLpy (Robitaille & Bressert, 2012), astropy (Astropy Collaboration et al., 2022), CMasher (van der Velden, 2020), matplotlib (Hunter, 2007), and numpy (van der Walt et al., 2011).References
- Akamatsu et al. (2017) Akamatsu, H., Fujita, Y., Akahori, T., et al. 2017, A&A, 606, A1
- Akamatsu et al. (2016) Akamatsu, H., Gu, L., Shimwell, T. W., et al. 2016, A&A, 593, L7
- Alvarez et al. (2022) Alvarez, G. E., Randall, S. W., Su, Y., et al. 2022, ApJ, 938, 51
- Astropy Collaboration et al. (2022) Astropy Collaboration, Price-Whelan, A. M., Lim, P. L., et al. 2022, ApJ, 935, 167
- Balboni et al. (2025) Balboni, M., Gastaldello, F., Bonafede, A., et al. 2025, arXiv e-prints, arXiv:2507.00133
- Bonafede et al. (2021) Bonafede, A., Brunetti, G., Vazza, F., et al. 2021, ApJ, 907, 32
- Bonjean et al. (2018) Bonjean, V., Aghanim, N., Salomé, P., Douspis, M., & Beelen, A. 2018, A&A, 609, A49
- Botteon et al. (2025) Botteon, A., Balboni, M., Bartalucci, I., Gastaldello, F., & van Weeren, R. J. 2025, A&A, 698, A55
- Botteon et al. (2020a) Botteon, A., Brunetti, G., van Weeren, R. J., et al. 2020a, ApJ, 897, 93
- Botteon et al. (2019a) Botteon, A., Cassano, R., Eckert, D., et al. 2019a, A&A, 630, A77
- Botteon et al. (2018) Botteon, A., Shimwell, T. W., Bonafede, A., et al. 2018, MNRAS, 478, 885
- Botteon et al. (2019b) Botteon, A., Shimwell, T. W., Bonafede, A., et al. 2019b, A&A, 622, A19
- Botteon et al. (2022) Botteon, A., Shimwell, T. W., Cassano, R., et al. 2022, A&A, 660, A78
- Botteon et al. (2020b) Botteon, A., van Weeren, R. J., Brunetti, G., et al. 2020b, MNRAS, 499, L11
- Botteon et al. (2024) Botteon, A., van Weeren, R. J., Eckert, D., et al. 2024, A&A, 690, A222
- Briggs (1995) Briggs, D. S. 1995, in American Astronomical Society Meeting Abstracts, Vol. 27, 1444
- Brüggen et al. (2021) Brüggen, M., Reiprich, T. H., Bulbul, E., et al. 2021, A&A, 647, A3
- Brunetti & Jones (2014) Brunetti, G. & Jones, T. W. 2014, Int. J. Mod. Phys. D, 23, 30007
- Brunetti & Vazza (2020) Brunetti, G. & Vazza, F. 2020, PRL, 124, 51101
- Bruno et al. (2023) Bruno, L., Brunetti, G., Botteon, A., et al. 2023, A&A, 672, A41
- Bulbul et al. (2016) Bulbul, E., Randall, S. W., Bayliss, M. B., et al. 2016, ApJ, 818, 131
- Caglar (2018) Caglar, T. 2018, MNRAS, 475, 2870
- Caglar & Hudaverdi (2017) Caglar, T. & Hudaverdi, M. 2017, MNRAS, 472, 2633
- Cassano et al. (2013) Cassano, R., Ettori, S., Brunetti, G., et al. 2013, ApJ, 777, 141
- Chandra et al. (2004) Chandra, P., Ray, A., & Bhatnagar, S. 2004, ApJ, 612, 974
- Churazov et al. (2021) Churazov, E., Khabibullin, I. I., Lyskova, N., Sunyaev, R., & Bykov, A. M. 2021, A&A, 651, A41
- Cuciti et al. (2021) Cuciti, V., Cassano, R., Brunetti, G., et al. 2021, A&A, 647, A51
- Cuciti et al. (2023) Cuciti, V., Cassano, R., Sereno, M., et al. 2023, A&A, 680, A30
- David & Kempner (2004) David, L. P. & Kempner, J. C. 2004, ApJ, 613, 831
- de Jong et al. (2022) de Jong, J. M., van Weeren, R. J., Botteon, A., et al. 2022, A&A, 668, A107
- Dey et al. (2019) Dey, A., Schlegel, D. J., Lang, D., et al. 2019, AJ, 157, 168
- Döner et al. (2025) Döner, S., Caglar, T., Smith, K. L., et al. 2025, ApJ submitted
- Donnert et al. (2013) Donnert, J. M., Dolag, K., Brunetti, G., & Cassano, R. 2013, MNRAS, 429, 3564
- Duchesne et al. (2024) Duchesne, S. W., Botteon, A., Koribalski, B. S., et al. 2024, PASA, 41, e026
- Duchesne et al. (2021a) Duchesne, S. W., Johnston-Hollitt, M., Offringa, A. R., et al. 2021a, PASA, 38, e010
- Duchesne et al. (2021b) Duchesne, S. W., Johnston-Hollitt, M., & Wilber, A. G. 2021b, PASA, 38, e031
- George et al. (2021) George, L. T., Kale, R., & Wadadekar, Y. 2021, Experimental Astronomy, 51, 235
- Govoni et al. (2019) Govoni, F., Orrù, E., Bonafede, A., et al. 2019, Science, 364, 981
- Gu et al. (2019) Gu, L., Akamatsu, H., Shimwell, T. W., et al. 2019, Nat. Ast., 3, 838
- Gupta et al. (2017) Gupta, Y., Ajit Kumar, B., Kale, H. S., et al. 2017, Current Science, 113, 707
- Hincks et al. (2022) Hincks, A. D., Radiconi, F., Romero, C. E., et al. 2022, MNRAS, 510, 3335
- Hoeft et al. (2021) Hoeft, M., Dumba, C., Drabent, A., et al. 2021, A&A, 654, A68
- Hu et al. (2025) Hu, D., Werner, N., Xu, H., et al. 2025, A&A, 694, A320
- Hunter (2007) Hunter, J. D. 2007, Comput. Sci. Eng., 9, 90
- Iapichino et al. (2008) Iapichino, L., Adamek, J., Schmidt, W., & Niemeyer, J. C. 2008, MNRAS, 388, 1079
- Intema et al. (2009) Intema, H. T., van der Tol, S., Cotton, W. D., et al. 2009, A&A, 501, 1185
- Johnston-Hollitt & Pratley (2017) Johnston-Hollitt, M. & Pratley, L. 2017, arXiv e-prints [1706.04930]
- Jonas (2009) Jonas, J. L. 2009, IEEE Proceedings, 97, 1522
- Kato et al. (2015) Kato, Y., Nakazawa, K., Gu, L., et al. 2015, PASJ, 67, 71
- Kaya et al. (2019) Kaya, H. I., Caglar, T., & Sert, H. 2019, MNRAS, 485, 4550
- Kim et al. (1989) Kim, K.-T., Kronberg, P. P., Giovannini, G., & Venturi, T. 1989, Nature, 341, 720
- Knowles et al. (2022) Knowles, K., Cotton, W. D., Rudnick, L., et al. 2022, A&A, 657, A56
- Machado et al. (2024) Machado, R. E. G., Volert, R. C., Albuquerque, R. P., Monteiro-Oliveira, R., & Lima Neto, G. B. 2024, ApJ, 970, 160
- Migkas et al. (2025) Migkas, K., Sommer, M. W., Schrabback, T., et al. 2025, A&A, 694, A45
- Mirakhor et al. (2022) Mirakhor, M. S., Walker, S. A., & Runge, J. 2022, MNRAS, 509, 1109
- Mirakhor et al. (2023) Mirakhor, M. S., Walker, S. A., & Runge, J. 2023, MNRAS, 522, 2105
- Murgia et al. (2024) Murgia, M., Govoni, F., Vacca, V., et al. 2024, MNRAS, 528, 6470
- Nunhokee et al. (2023) Nunhokee, C. D., Bernardi, G., Manti, S., et al. 2023, MNRAS, 522, 4421
- Offringa (2016) Offringa, A. R. 2016, A&A, 595, A99
- Offringa et al. (2010) Offringa, A. R., de Bruyn, A. G., Biehl, M., et al. 2010, MNRAS, 405, 155
- Offringa et al. (2014) Offringa, A. R., McKinley, B., Hurley-Walker, N., et al. 2014, MNRAS, 444, 606
- Offringa & Smirnov (2017) Offringa, A. R. & Smirnov, O. M. 2017, MNRAS, 471, 301
- Offringa et al. (2012) Offringa, A. R., van de Gronde, J. J., & Roerdink, J. B. 2012, A&A, 539, A95
- Omiya et al. (2023) Omiya, Y., Nakazawa, K., Matsushita, K., et al. 2023, PASJ, 75, 37
- Pignataro et al. (2024a) Pignataro, G., Bonafede, A., Bernardi, G., et al. 2024a, A&A, 691, A99
- Pignataro et al. (2024b) Pignataro, G., Bonafede, A., Bernardi, G., et al. 2024b, A&A, 685, L10
- Pignataro et al. (2024c) Pignataro, G., Bonafede, A., Bernardi, G., et al. 2024c, A&A, 682, A105
- Planck Collaboration XXII (2016) Planck Collaboration XXII. 2016, A&A, 594, A22
- Rajpurohit et al. (2025) Rajpurohit, K., Botteon, A., O’Sullivan, E., et al. 2025, ApJ, 992, 88
- Reddy et al. (2017) Reddy, S. H., Kudale, S., Gokhale, U., et al. 2017, Journal of Astronomical Instrumentation, 6, 1641011
- Robitaille & Bressert (2012) Robitaille, T. P. & Bressert, E. 2012, APLpy: Astronomical Plotting Library in Python
- Roettiger et al. (1997) Roettiger, K., Loken, C., & Burns, J. O. 1997, ApJ, Supplement, 109, 307
- Roettiger et al. (1999) Roettiger, K., Stone, J. M., & Burns, J. O. 1999, ApJ, 518, 594
- Roy et al. (2010) Roy, J., Gupta, Y., Pen, U.-L., et al. 2010, Experimental Astronomy, 28, 25
- Sakelliou & Ponman (2004) Sakelliou, I. & Ponman, T. J. 2004, MNRAS, 351, 1439
- Sikhosana et al. (2025) Sikhosana, S. P., Hilton, M., Bernardi, G., et al. 2025, A&A, 698, L17
- Stuardi et al. (2025) Stuardi, C., Botteon, A., Sereno, M., et al. 2025, A&A, 695, L16
- van der Velden (2020) van der Velden, E. 2020, J. Open Source Softw., 5, 2004
- van der Walt et al. (2011) van der Walt, S. J., Colbert, S. C., & Varoquaux, G. 2011, Comput. Sci. Eng., 13, 22
- van Diepen et al. (2018) van Diepen, G., Dijkema, T. J., & Offringa, A. R. 2018, DPPP: Default Pre-Processing Pipeline
- van Weeren et al. (2019) van Weeren, R. J., de Gasperin, F., Akamatsu, H., et al. 2019, Space Sci. Rev., 215, 16
- van Weeren et al. (2025) van Weeren, R. J., Osinga, E., Brunetti, G., et al. 2025, MNRAS submitted
- van Weeren et al. (2021) van Weeren, R. J., Shimwell, T. W., Botteon, A., et al. 2021, A&A, 651, A115
- Vazza et al. (2019) Vazza, F., Ettori, S., Roncarelli, M., et al. 2019, A&A, 627, A5
- Vazza et al. (2017) Vazza, F., Jones, T. W., Brüggen, M., et al. 2017, MNRAS, 464, 210
- Venturi et al. (2022) Venturi, T., Giacintucci, S., Merluzzi, P., et al. 2022, A&A, 660, A81
- Werner et al. (2008) Werner, N., Finoguenov, A., Kaastra, J. S., et al. 2008, A&A, 482, L29
- XRISM Collaboration et al. (2025a) XRISM Collaboration, Audard, M., Awaki, H., et al. 2025a, PASJ in press
- XRISM Collaboration et al. (2025b) XRISM Collaboration, Audard, M., Awaki, H., et al. 2025b, ApJ, Letters, 982, L5
- XRISM Collaboration et al. (2025c) XRISM Collaboration, Audard, M., Awaki, H., et al. 2025c, ApJ, Letters, 985, L20
- XRISM Collaboration et al. (2025d) XRISM Collaboration, Audard, M., Awaki, H., et al. 2025d, PASJ, 77, S242
- XRISM Collaboration et al. (2025e) XRISM Collaboration, Audard, M., Awaki, H., et al. 2025e, Nature, 638, 365
- ZuHone et al. (2018) ZuHone, J. A., Kowalik, K., Öhman, E., Lau, E. T., & Nagai, D. 2018, ApJ, Supplement, 234, 4